# A practical identifiability criterion leveraging weak-form parameter estimation

### v2

@@ 6 Acknowledgments
Research reported in this publication was supported in part by the NIGMS Division of Biophysics, Biomedical Technology and Computational Biosciences grant R35GM149335, in part by the NSF Division Of Environmental Biology (EEID grant DEB-2109774), and in part by the NIFA Biological Sciences (grant 2019-67014-29919).
@@ 7 Data and Code Availability
Appendix A Weak form Estimation of Nonlinear Dynamics (WENDy)
Recently, we have developed the Weak form Estimation of Nonlinear Dynamics (WENDy) method for parameter estimation method [26], which we will briefly describe here. We begin by considering the $i$th variable $u_{i}$ in a $d$-dimensional ordinary differential equation (ODE)
$\dot{u}_{i}=\Theta(u)\mathbf{w}_{i}$ (14)
with row vector of $d$ solution states
$u(t;\mathbf{w}):=(u_{1}(t;\mathbf{w}),\dots,u_{d}(t;\mathbf{w}))$
and row vector of $J$ features $\Theta(u):=(f_{1}(u),\dots,f_{J}(u))$. Each feature $f_{j}:\mathbb{R}^{d}\to\mathbb{R}$ is assumed to be $C^{2}$. We seek to identify the vector of unknown parameters $\mathbf{w}=[\mathbf{w}_{1}^{T}|\cdots|\mathbf{w}_{d}^{T}]^{T}\in\mathbb{R}^{Jd}$ from observations of $u$.
We let the superscript $\star$ notation denote quantities based on the true (noise-free and unknown) parameter values ($\mathbf{w}^{\star}$) and let $u^{\star}:=u(t;\mathbf{w}^{\star})$ and $\Theta^{\star}:=\Theta(u^{\star})$. We assume that at each timepoint $t\in\mathbf{t}=(t_{0},\dots,t_{M})$ measurements of the system are observed with additive noise
$U(t)=u^{\star}(t)+\varepsilon(t)$ (15)
where each element of $\varepsilon(t)$ is i.i.d. $\mathcal{N}(0,\sigma^{2})$. Using bold-face variables to represent evaluation at the timegrid $\mathbf{t}$, our observations then consist of samples $\mathbf{U}:=\mathbf{u}^{\star}+\boldsymbol{\varepsilon}\in\mathbb{R}^{(M+1)\times d}$.
To arrive at a weak-form framework to estimate $\mathbf{w}^{\star}$ from $\mathbf{U}$, we consider a $C^{\infty}$ bump function $\phi$ for which the function and all derivatives are compactly supported in the time interval $[0,T]$ (e.g. $\phi\in C_{c}^{\infty}([0,T])$). Multiplying both sides of (14) by $\phi$, and integrating by parts over 00 to $T$, we obtain
$-\int_{0}^{T}\dot{\phi}u\textsf{d}t=\int_{0}^{T}\phi\Theta(u)\textsf{mat}(\mathbf{w})\textsf{d}t,$ (16)
where “$\mathsf{mat}$” is the matricization operation. Using a set of these test function $\{\phi_{k}\}_{i=1}^{K}$, (16) can be discretized at observed data $\mathbf{U}$, yielding approximated integrals $-\dot{\Phi}_{k}\mathbf{U}=\Phi_{k}\Theta(\mathbf{U})W$ and corresponding linear system $\mathbf{G}\mathbf{w}^{\star}\approx\mathbf{b}$, where
$\mathbf{G}$ $:=[\mathbb{I}_{d}\otimes(\Phi\Theta(\mathbf{U}))],$
$\mathbf{b}$ $:=-\mathsf{vec}(\dot{\Phi}\mathbf{U}).$
A simple ordinary least-squares framework for identifying $\mathbf{w}^{\star}$ does not address the Errors-In-Variables nature of the problem (given the dependence of both $\mathbf{G}$ and $\mathbf{b}$ on the noisy data). To address this, the WENDy algorithm uses iterative reweighting to approximate the correct underlying covariance structure. For this purpose, consider the residual
$\mathbf{r}(\mathbf{U},\mathbf{w}):=\mathbf{G}\mathbf{w}-\mathbf{b},$ (17)
which can be decomposed into several components
$\mathbf{r}(\mathbf{U},\mathbf{w})=\underbrace{(\mathbf{G}-\mathbf{G}^{\star})\mathbf{w}}_{\begin{array}[]{c}\mathbf{e}_{\Theta}\end{array}}+\underbrace{\mathbf{G}^{\star}(\mathbf{w}-\mathbf{w}^{\star})}_{\begin{array}[]{c}\mathbf{r}_{0}\end{array}}+\underbrace{(\mathbf{G}^{\star}\mathbf{w}^{\star}-\mathbf{b}^{\star})}_{\begin{array}[]{c}\mathbf{e}_{\text{int}}\end{array}}-\mathbf{b}^{\boldsymbol{\varepsilon}}.$ (21)
Here, $\mathbf{r}_{0}$ is the residual without measurement noise or integration errors, and $\mathbf{e}_{\text{int}}$ is the numerical integration error induced by the quadrature. The leftover terms $\mathbf{e}_{\Theta}-\mathbf{b}^{\boldsymbol{\varepsilon}}$ can be Taylor expanded around the data $\mathbf{U}$,
$\mathbf{e}_{\Theta}-\mathbf{b}^{\boldsymbol{\varepsilon}}=\mathbf{L}_{\mathbf{w}}\mathsf{vec}(\boldsymbol{\varepsilon})+\mathbf{h}(\mathbf{U},\mathbf{w},\boldsymbol{\varepsilon})$ (22)
where $\mathbf{h}(\mathbf{U},\mathbf{w},\boldsymbol{\varepsilon})$ denotes higher-order terms in the measurement errors $\boldsymbol{\varepsilon}$. The first order matrix in the expansion (22) is
$\mathbf{L}_{\mathbf{w}}:=[\mathsf{mat}(\mathbf{w})^{T}\otimes\Phi]\nabla\Theta\mathbf{K}+[\mathbb{I}_{d}\otimes\dot{\Phi}],$
where $\mathbf{K}$ is the commutation matrix such that $\mathbf{K}\textsf{vec}(\boldsymbol{\varepsilon})=\textsf{vec}(\boldsymbol{\varepsilon}^{T})$. The matrix $\nabla\Theta$ contains derivatives of the features (for details see [26]). Under the Gaussian noise assumption, when $\mathbf{w}=\mathbf{w}^{\star}$ and the integration error is negligible, to first order we have
$\mathbf{G}\mathbf{w}^{\star}-\mathbf{b}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{L}_{\mathbf{w}^{\star}}(\mathbf{L}_{\mathbf{w}^{\star}})^{T}).$ (23)
This leads to an iterative alrogithm, where the covariance matrix $\mathbf{C}^{(n)}$ is updated using the weights $\mathbf{w}^{(n)}$ step $n$, for regularization parameter $\alpha$,
$\mathbf{L}^{(n)}$ $\leftarrow[\textsf{mat}(\mathbf{w}^{(n)})^{T}\otimes\Phi]\nabla\Theta(\mathbf{U})\mathbf{K}+[\mathbb{I}_{d}\otimes\dot{\Phi}]$
$\mathbf{C}^{(n)}$ $=(1-\alpha)\mathbf{L}^{(n)}(\mathbf{L}^{(n)})^{T}+\alpha\mathbf{I}$
and the weights are subsequently updated as
$\mathbf{w}^{(n+1)}$ $\leftarrow(\mathbf{G}^{T}(\mathbf{C}^{(n)})^{-1}\mathbf{G})^{-1}\mathbf{G}^{T}(\mathbf{C}^{(n)})^{-1}\mathbf{b}.$
@@ Appendix A Weak form Estimation of Nonlinear Dynamics (WENDy)
Appendix B Derivations of weak-form input-ouput equations for SIR model
Consider model equation, which is presented again below,
@@ Appendix B Derivations of weak-form input-ouput equations for SIR model
From the $\dot{S}$ equation we can also obtain equation the following by taking the derivative,
@@
From the $\dot{S}$ and $\dot{I}$ equations we can obtain the following equation by taking the derivative,
@@
Next we substitute equation (29) in equation (28) to obtain equation (10), an input-output equation for the SIR model for observed state variable $I,$ which we present again below,
@@
Notice that three of the terms on the right hand side of (10) can be written in the form (1), or in other words can be written a derivative of a function of only $I$. However, the term $\frac{\dot{I}^{2}}{I}$ cannot be put in this form. We re-write equation (10) in the form (3) by converting as many terms as possible to form (1) to obtain equation the following,
@@
So, using equation (27) we can then re-write the term which is not integrable by parts as follows,
@@
Let the total population $N=S(t)+I(t)+R(t)$ be constant. Then, $S(t)=N-I(t)-R(t)$. From the equation (9) we know,
$R(t)=\int_{0}^{t}\alpha I(s)\textsf{d}s.$
Then from our assumption of constant population we are able to generate the following equation from equation (32), which is given again below,
@@
Finally, we can write equation (13), a weak-form SIR model that reveals parameter $\beta$ using only observations of state $I$, which is given again below,

### v3

@@ 6 Acknowledgments
Research reported in this publication was supported in part by the NIGMS Division of Biophysics, Biomedical Technology and Computational Biosciences grant R35GM149335, in part by the NSF Division Of Environmental Biology (EEID grant DEB-2109774), and in part by the NIFA Biological Sciences (grant 2019-67014-29919). We thank Rainey Lyons (University of Colorado) for insightful discussions on test function formulation. We also thank the reviewers for feedback on the input-output definition of structural identifiability.
@@ 7 Data and Code Availability
Appendix A Test Function Selection
It is well known that the choice of test function has a significant impact on the performance of weak-form methods. There have been several efforts to wisely select test functions, including spectral matching [26] and quadrature error minimization [28]. And, more recently, a method has been proposed to construct test functions for accurate parameter estimation [55]. However, optimal selection of test functions remains an open problem. In this work, we choose the test function optimally, based on knowing the true parameter. This ensures that the focus of this work can remain on the impact of constructing weak-form Input/Output equations.
We select the degree of the polynomial test function and the radius of compact support for each example presented in this paper by observing the relative error in parameter estimates. For model (9), we select our test functions to be $12^{\text{th}}$ order polynomials with a radius of support $a=0.52$. Figure 9 shows the relative error in $w$ for polynomial test function choices of order 6 through 18. We observed a significant improvement in accuracy when moving to a $12^{\text{th}}$ order polynomial, but did not gain accuracy by increasing the order further. Figure 10 shows the relative error in $w$ for compact support for radii from $a=[0.45,0.65]$.
Figure 9: The relative error in parameters $w$ when fitting model (9) to 400 observations using WENDy for polynomial test function choices of degree =[6,18]=[6,18]. The red star indicates the $12^{\text{th}}$ order polynomial used in this study.
Figure 10: The relative error in parameters $w$ when fitting model (9) to 400 observations using WENDy for polynomial test function radii of support $a=[0.45,0.65]$. The red star indicates the choice of $a=0.52$ used in this study.
For model (17), we select our test functions to be $20^{\text{th}}$ order polynomials with a radius of support $a=11$. Figure 11 shows the relative error in $\beta$ for polynomial test functions of degree =[10,26]=[10,26]. Figure 12 shows the relative error in $\beta$ for compact support for radii from $a=[5,15]$.
Figure 11: The relative error in parameters $\beta$ when fitting model (17) to 31 observations using WENDy for polynomial test function choices of degree =[10,26]=[10,26]. The red star indicates the $20^{\text{th}}$ order polynomial used in this study.
Figure 12: The relative error in parameters $\beta$ when fitting model (9) to 31 observations using WENDy for polynomial test function radii of support $a=[5,15]$. The red star indicates the choice of $a=11$ used in this study.
@@ Appendix A Test Function Selection
Appendix B Derivations of weak-form input-output equations for SIR model
Consider the model equation, which is presented again below,
@@ Appendix B Derivations of weak-form input-output equations for SIR model
From the $\dot{S}$ equation, we can also obtain the following equation by taking the derivative,
@@
From the $\dot{S}$ and $\dot{I}$ equations, we can obtain the following equation by taking the derivative,
@@
Next, we substitute equation (23) in equation (22) to obtain equation (14), an input-output equation for the SIR model for observed state variable $I,$ which we present again below,
@@
Notice that three of the terms on the right-hand side of (14) can be written in the form (1), or in other words, can be written a derivative of a function of only $I$. However, the term $\frac{\dot{I}^{2}}{I}$ cannot be put in this form. We re-write equation (14) in the form (3) by converting as many terms as possible to form (1) to obtain equation the following,
@@
Thus, using equation (21), we can then rewrite the term which is not integrable by parts as follows,
@@
Note that on the right-hand side of equation (26) is written in the terms of the state variables, $S$ and $I$, and the test function and its derivative, $\phi$ and $\dot{\phi}$, and does not include derivatives of the state variables (as desired). We assume to only observe the state variable, $I$, and therefore, we cannot use equation (26) directly. From the equation (13) we know,
$\begin{split}\frac{dR}{dt}=\alpha I\Rightarrow R(t)=\int_{0}^{t}\alpha I(s)\textsf{d}s.\end{split}$
Note, here we assume a naive population, i.e., $R(0)=0$. Given that we observe the infected compartment, we are able to compute $R(t)$ through numeric integration.
Let the total population $N=S(t)+I(t)+R(t)$ be constant and known. Then, S(t)=NI(t)R(t)S(t)=N-I(t)-R(t) can be computed from only observations of the state variable $I(t)$. Then, using our assumption of a known constant population, we are able to generate the following equation from equation (26), which is given again below,
@@
Finally, we can write equation (17), a weak-form SIR model that reveals parameter $\beta$ using only observations of state $I$ and a known population size, which is given again below,
@@
Appendix C Structural Identifiability of Example Models
Example 1
Below, we present again equation (8), the input-output equation for model 7 in the strong form,
$\begin{split}\ddot{x}_{1}(t)x_{1}(t)^{2}+2\ddot{x}_{1}(t)x_{1}(t)+\ddot{x}_{1}(t)+w_{1}\left(x_{1}(t)^{2}+x_{1}(t)\right)\\ +w_{2}\left(\dot{x}_{1}(t)x_{1}(t)^{2}+2\dot{x}_{1}(t)x_{1}(t)\right)+w_{3}\dot{x}_{1}(t)=0,\end{split}$
where $w_{1}=k_{21}V_{e}$, $w_{2}=k_{12}+k_{21}$, and $w_{3}=k_{12}+k_{21}+V_{e}$. In order to verify the structural identifiability from this input-output representation, we can consider Definition 2. However, for this to be a sufficient condition for the structural identifiability, the rank of the Wronskian of the differential monomials of the input-output representation must be checked [39]. Recently, [19] have developed a projection-based approach analogous to the differential elimination approach that takes into consideration conditions for the equivalence of multi and single experiment identifiability. We confirm both the form of equation (8) and that the model parameters are globally structurally identifiable using find_ioequations and assess_identifiability commands from the StructuralIdentifiability package in Julia [19]. While StructuralIdentifiability.jl also produces the above strong form input-output equation, we chose to use the Differential Algebra v4 package to generate equation (8) in order to perform the integration by parts procedure that is needed to generate the weak-form in this study.
While we cannot directly test the identifiability of the weak-form input-output equation (9), we can confirm the unique recoverablity of the parameters in the discretized WENDy equation $\mathbf{G}\mathbf{w}^{*}=\mathbf{b}$. In the case of no noise,
$\mathbf{G}=\begin{bmatrix}\mathbf{\Phi}\frac{\mathbf{y}}{\mathbf{y}+1}&-\mathbf{\dot{\Phi}}\frac{\mathbf{y}^{2}}{\mathbf{y}+1}&\mathbf{\dot{\Phi}}\frac{1}{\mathbf{y}+1}\end{bmatrix}$
For the chosen test function matrix, 400 noise-free observations, and assuming model parameters $k_{12}=5,$ $k_{21}=1,$ and $V_{e}=6$, we verify G is full rank using the SVD algorithm in Matlab with a tolerance of 1e121e-12. This result supports that, at least local to the parameter choices used in this example, the transformation from the strong to weak-form input-output system did not result in a loss of parameter identifiability. We also verify that $\text{rank}(\mathbf{\Phi})=\text{rank}(\dot{\mathbf{\Phi}})=K$ with a tolerance of 1e121e-12, for this example $K=15$. To further support this claim, we graph the terms $C_{1},$ $C_{2}$, and $C_{3}$ from the weak-form input-output equation (9) in Figure 13. Notice that the convolved terms each provide a unique signal.
Figure 13: Dynamics of terms $C_{1},$ $C_{2}$, and $C_{3}$ from the weak-form input-output equation (9) for parameters $k_{12}=5,$ $k_{21}=1,$ and $V_{e}=6$.
Example 2
Similarly, for the SIR example, we confirm both the form of equation (14) and that the model parameters are globally structurally identifiable using find_ioequations and assess_identifiability commands from the StructuralIdentifiability package in Julia [19].
While we cannot directly test the identifiability of the weak-form input-output equation (17), we can confirm the unique recoverability of the parameters in the discretized WENDy equation $\mathbf{G}\mathbf{w}^{*}=\mathbf{b}$. In the case of no noise,
$\mathbf{G}=\begin{bmatrix}\mathbf{\Phi}\left(\mathbf{y}^{2}-(N-\mathbf{R}-\mathbf{y})^{2}\right)&\mathbf{\dot{\Phi}}\left(\mathbf{y}^{2}+2(N-\mathbf{R}-\mathbf{y})\mathbf{y}\right)\end{bmatrix}$
For the chosen test function matrix, 31 noise-free observations, and assuming model parameters N=10,000, $S_{0}=N-1$, $I_{0}=1$, $R_{0}=0$, $\beta=\frac{5.5}{N}$, and $\gamma=5$, we verify G is full rank using the SVD algorithm in Matlab with a tolerance of 1e121e-12. This result supports that, at least local to the parameter choices used in this example, the transformation from the strong to weak-form input-output system did not result in a loss of parameter identifiability. We also verify that $\text{rank}(\mathbf{\Phi})=\text{rank}(\dot{\mathbf{\Phi}})=K$ with a tolerance of 1e121e-12, for this example $K=4$.
Appendix D Example 1 $(e,q)$-identifiability with lognormal noise
We estimate the parameters $w_{1}$, $w_{2}$, and $w_{3}$ using WENDy for simulated data with log-normal multiplicative observation error. We vary the multiplicative observation error ratio in the simulated data between $e=[0\%,20\%]$ such that the standard deviation of the error is $\sigma=e\log(\text{RMS}(\Omega(t)))$, and at each multiplicative observation error ratio, we generate 1,000 noisy datasets. The (e,q)-identifiability for $e=[0\%,20\%]$ and $q=[1\%,100\%]$ of model (9) is given in Figure 14A. Unlike in the case of additive noise, we find that the weak form model is (10,20)-identifiable, or generally practically identifiable.
Figure 14: Using WENDy with equations (9) we find the SIR model to be generally practically identifiable for multiplicative noise between $0\%,20\%]$. (A.) The area in blue denotes where the model is (e,q)-identifiable, and the area in white denotes where the model is not (e,q)-identifiable. The (e,q)-identifiability was determined from 1,000 samples at each respective error level, e. The red star marks the $(10,20)$-identifiability criterion. (B.) The relative error determined from 1,000 samples at each respective error level, e. (C.) The proportion of estimated $95\%$ confidence intervals at $e=[0\%,20\%]\%$ that contain the true value of $w_{i}$.
We additionally compare the (e,q)-identifiability of model equations (9) to the average relative error and the coverage of the $95\%$ confidence intervals generated at each additive error ratio. The relative error generated for 1,000 simulated datasets at each multiplicative observation error level is given in Figure 14B, and the coverage of the $95\%$ confidence intervals for $w_{1}$, $w_{2}$, and $w_{3}$ is given in Figure 14C. Here, the (e,q)-identifiability and the relative error criteria are in agreement.